OpenStudy (anonymous):

If f(x)=sinx^2, find f'(x) and hence find integrate(xcosx^2)dx okay i think i'm having a bit of problem with the front part.. is it the same as sin^2x??

OpenStudy (kainui):

So is this your integral? \[\LARGE \int\limits x \cos (x^2)dx\] I'm just making sure the squared part is there.

OpenStudy (anonymous):

wait dang the prime means to differentiate doesn't it?

OpenStudy (anonymous):

oh yes yup

OpenStudy (anonymous):

does that make f'(x)=cos2x?

OpenStudy (kainui):

Almost, do know what the chain rule is? Try to use the definition to get the right answer

OpenStudy (anonymous):

it isn't cos2x? D:

OpenStudy (kainui):

So for example if I have \[\LARGE f(x)=\sin(x^3)\] then the derivative is \[\LARGE f'(x)=\cos(x^3)*3x^2\] since I took the derivative of the outside and multiplied it by the derivative of the inside.

OpenStudy (anonymous):

2sincosx^2?